OpenStudy (anonymous):

Solve the surface integral: \[\int\limits_{}^{}\int\limits_{S}^{}xz \ dS\] S is the part of the plane 2x+2y+z=4 that lies in the first octant

OpenStudy (anonymous):

I don't even know where to begin with this. I think you have to find r_u and r_v to start but I don't know how to find those. @TuringTest @UnkleRhaukus @eseidl @experimentX @phi

OpenStudy (anonymous):

Ok so I found the first step, but I'm not sure how they got those points or those vectors

OpenStudy (anonymous):

First you have to parametrize the surface in terms of x and y in this case. you can solve the equation of the plane for z and that puts the plane in terms of x and y. so we get the parametrized surface in vector format to be: \[\Phi(x,y)=\left( x,y, g(x,y) \right) \] which becomes: \[\Phi(x,y)=\left( x, y, 4-2x-2y \right)\]

OpenStudy (anonymous):

Then you have to find the tangent vector to the parametrized surface and take the cross product of those to get the normal vector.

OpenStudy (anonymous):

I got: \[T _{x}=(1,0,-2) \] \[T _{y}=(0, 1, -2)\]

OpenStudy (anonymous):

This seems a lot more familiar, but it doesn't match up with what the answer key showed?

OpenStudy (anonymous):

The goal is to get \[\int\limits_{?}^{?}\int\limits_{?}^{?}_{D} f(x, y, g(x,y)\left| n(x,y) \right|\]

OpenStudy (anonymous):

Would be it: \[\int\limits\limits_{?}^{?}\int\limits\limits_{?}^{?}x*(4-2x-2y)*3 \ dS\]

OpenStudy (anonymous):

I get the normal vector to be \[n=(2, 2, 1)\] then taking its magnitude I get: \[\sqrt{2^{2}+2^{2}+1^{2}}=\sqrt{5}\]

OpenStudy (anonymous):

Yeah, but you found r_x = (1,0,-2) and r_y=(0,1,-2) so the magnitude of the cross product would be |2i+2j+1k|=sqrt(9)=3 Where did n come from?

OpenStudy (anonymous):

Nevermind your n is the cross product, the only mistake is that the magnitude is 3

OpenStudy (anonymous):

yes true.....sorry bad arithmetic

OpenStudy (anonymous):

I'm confused as to what comes next

OpenStudy (anonymous):

ok i think it should be \[\int\limits_{0}^{2}\int\limits_{0}^{2} 3x(4-2x-2y) dxdy \] bear with me, I just learned these a couple weeks ago and maybe a bit shaky.....

OpenStudy (anonymous):

I completely understand. I can't stand these! Ok, I got the same thing you did for the function, but how did you find the bounds?

OpenStudy (anonymous):

yeah they suck!!! I have a test coming up in this soon too.... Since we are finding they told us the surface is in the first octant I set y and z = 0 and solved for x and then did the same for y. If you draw out the plane you will see where the plane intersects the x and y axis is 2

OpenStudy (anonymous):

Got it! Thank you! You have no idea how long I've been stuck on this problem!!

OpenStudy (anonymous):

did we get the correct answer using that method?

OpenStudy (anonymous):

I'll check with WA

OpenStudy (anonymous):

cool let me know, I hope I didn't lead you astray, and I have the complete idea how long you've been stuck!!! trust me....I'm in the same boat! Calc 3 is a bear!

OpenStudy (anonymous):

Ok, so wolfram alpha says it is -8 and the answer from the site I found says the answer is 20, but someone commented saying that they found the answer as 4, so I'm not sure what to think. Everything looks correct though so I think I'm going to stick with -8! Thanks again and best of luck on your test!

OpenStudy (anonymous):

I also got -8 for the answer. Good luck with your class!!!!

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

This may not be of interest, but here is a write-up on doing this problem